Two masses m₁ = 350 kg and m₂ = 350 kg are at a distance of 22 m from each other. Find the magnitude and direction of the gravitational field at point A. Data: h = 9 m, G = 6.674×10–11 N·m²/kg². 


1
Expert's answer
2021-03-23T08:14:07-0400
"E=(6.67\cdot10^{-11})\frac{350}{9^2}-(6.67\cdot10^{-11})\frac{350}{(22-9)^2}\\
E=1.5\cdot10^{-10}\frac{N}{kg}"

Direction: towards first mass.